OpenStudy (anonymous):

find formulas f+g,f-g,fg n f/g n state the domain f(x)=2x , g(x)=x^2+1

OpenStudy (anonymous):

f+g = x^2 +2x +1 f-g = -x^2 +2x-1 fg = 2x^3 +2x f/g = 2x/x^2 +1 Domain is all R for all the functions